OpenStudy (austinl):

Differential Equations Initial Value Problem I worked it out, but I arrived at a different answer than is given in my text book. \(ty\prime + (t+1)y=t\) \(y(\ln(2))=1,~t>0\)

OpenStudy (austinl):

I arrived at y = t + 1 - ln(2) My books answer.... \(y=\dfrac{(t-1+2e^{-t})}{t},~t\ne0\)

terenzreignz (terenzreignz):

You have a new DE and you didn't invite me? T.T

OpenStudy (austinl):

@terenzreignz :P

OpenStudy (austinl):

My way is clearly incorrect. I think I need to treat the t+1 term differently. Any thoughts?

terenzreignz (terenzreignz):

I say go for that integrating factor again... After dividing everything by t, then we get the integrating factor ro be \[\Large \mu = \int \frac{t+1}{t}dt\]

OpenStudy (austinl):

Darn it.... stupid stupid stupid stupid stupid Austin.

terenzreignz (terenzreignz):

Whoops... I meant...\[\LARGE e^\mu\] as the integrating factor...

OpenStudy (austinl):

\(\huge{\mu(t)=te^t}\)

OpenStudy (austinl):

Look ok?

terenzreignz (terenzreignz):

That's the integrating factor itself, so... okay, great :D Multiply everything with it... you get \[\Large te^t y' + (t+1)e^ty = te^t\]

terenzreignz (terenzreignz):

That's the integrating factor itself, so... okay, great :D Multiply everything with it... you get \[\Large te^t y' + (t+1)e^ty = te^t\]

OpenStudy (austinl):

\(\dfrac{(t+1)}{t}\) Right?

terenzreignz (terenzreignz):

What do you mean? THIS term? \[\Large te^t y' +\color{red}{ (t+1)e^ty} = te^t\]

OpenStudy (austinl):

Yes.

OpenStudy (anonymous):

the standard equation for a linear differential equation is y' +Py = Q the first step towards solving the differential equation lies in identifying weather the given equation is in standard form or not. This can be achieved by comparing the given equation with the standard form. By doing so, we find that the given eq. isn't in the standard form so we must do certain operations to bring it to the standard form in this case We divide the both sides with 't' to obtain y' + (t+1)/t = 1 here, P=(t+1)/t or P =t/t +1/t or P= 1+1/t IF = e^int(p dt) IF =e^[int(1) + int(1/t )] IF =e^[t+log t] IF = e^t + e^log t IF = e^t + t the solution is given by ye^int(p dx) = int(Q e^int(p dx))

terenzreignz (terenzreignz):

Remember, we started with this: \[\Large t y' + (t+1)y = t\] divided everything by t \[\Large y' + \frac{(t+1)}ty =1\] Multiplied everything by \(\large te^t\), the integragint factor, arriving finally at: \[\Large te^t y' + (t+1)e^ty = te^t\] Questions? :P

OpenStudy (austinl):

Oh... duh Austin. I swear I don't function without caffeine. I feel like I just went full dipstick.

terenzreignz (terenzreignz):

Okay, well, I don't need caffeine, but some mocha ice cream would be awsum ^_^ Anyway, left side is now \[\Large \frac{d}{dt}[te^ty]= te^t\] Let's have at it Austin :3

OpenStudy (austinl):

Yep, I have worked that out. And then I have arrived at C=2

OpenStudy (austinl):

\(te^ty=(t-1)e^t+2\)

terenzreignz (terenzreignz):

Does that work out?

terenzreignz (terenzreignz):

Oh, it IS right, then?

OpenStudy (austinl):

I believe so!

terenzreignz (terenzreignz):

Then... ice cream time :D More ice cream, that is ^_^

terenzreignz (terenzreignz):

Lovely :D Well, I'm off ^_^ Have fun XD ------------------------------ Terence out

OpenStudy (austinl):

Thanks for all of your help Terence!